Ellie is 33 years old. This is three times as old as Sophie was when Ellie was as old as Sophie is now."
How old is Sophie now?

Answer by scott8148(6628)

let x="Sophie now"

NOW __ Ellie is 33, Sophie is x

THEN ___ Ellie is x, Sophie is 11

x-11=33-x __ 2x=44